Solve the following systems by substitution. Substitute the expression in the second equation into the first equation and solve. (If the system has no solution, enter NONE for each answer.)
x-y= -4
y= 3x+6

Answer by robertb(5830)

Replace y in the 1st equation by 3x+6 (from the bottom equation):
x - (3x+6) = -4,
x - 3x-6 = -4,
-2x = 2,
x = -1. Then y = 3*-1 + 6 = -3+6 = 3.